# An Open-Source Soft Robotic Platform for Autonomous Aerial Manipulation in the Wild

### v1

@@ Abstract
Figure 1: Our autonomous aerial platform demonstrates its aerial manipulation capabilities with onboard only perception. Left: Lifting a water bottle indoors after successful detection and grasp planning. Right: Grasping a plush toy outdoors.
@@ 1 Introduction
Aerial manipulation, which combines the agility of aerial vehicles with the functional capabilities of manipulators, poses a significant challenge in robotics. Traditional mobile manipulation platforms, typically composed of a mobile base and a manipulator, are designed to interact with their environment and perform various tasks. Aerial manipulators, using multicopters as their mobile base, offer significant advantages over ground-based platforms such as legged robots [1] and wheeled robots [2], particularly in scenarios requiring high traversal speeds or navigation through impassable terrain, such as during floods or in debris-strewn areas.
In recent years, aerial manipulation has seen significant advances using soft robotic manipulators [3, 4, 5, 6]. The compliant properties of soft manipulators relax the otherwise high constraints on precisely positioning the manipulator when grasping objects. Furthermore, soft manipulators also dampen the impact on the dynamics of the flying platform, which eases the burden on flight controllers during each grasp.
However, a significant limitation of existing aerial manipulation platforms is their dependence on external perception systems, such as motion capture systems [4, 5, 7], fiducial markers [3, 8], and VIVE trackers [9], to perform self-localization and target object localization. These systems simplify the research problem by providing nearly perfect pose information, but also present two core issues.
@@
To overcome these limitations, we propose a new platform that relies exclusively on onboard perception sensors for aerial manipulation. For self-localization, we use an existing simultaneous localization and mapping (SLAM) system [10], allowing the platform to estimate its state using visual observations alone. We implement a learning-based pipeline to localize target objects through RGB-D segmentation and point cloud analysis to determine target grasp locations.
@@ 2 Related Work
The integration of aerial systems with grasping mechanisms has gained considerable attention in recent years [11]. Research has demonstrated the potential of aerial grasping using soft grippers, which adapt to the shapes of objects due to their passive compliance [3]. Further advances include rapid grasping platforms [4] and efforts to achieve more autonomous grasping without relying on fiducial markers [12].
However, many existing systems still depend on motion capture systems to provide the precise location of the aerial platform or the object. For example, several works have focused on improving flying manipulators with stiffness control to improve object interaction dynamics [13] and developing lightweight, low activation force grippers designed specifically for aerial grasping applications [14]. In addition, Wu et. al. [15] has investigated optimizing quadcopter designs to enhance aerial grasping performance.
Despite these advancements, the reliance on motion capture systems for self-localization and object priors for target localization remains a common theme. More recent work has demonstrated successful aggressive aerial grasping using onboard perception by employing visual-inertial odometry for self-localization and CAD models for object recognition [5]. Our approach differs from this by leveraging a learning-based target localization system for zero-shot video segmentation integrated with a SLAM pipeline. This method eliminates the need for prior assumptions regarding object geometry or the environment, enhancing the flexibility and applicability of aerial grasping systems in unstructured settings.
@@ 3.1.1 Platform
[…]
A custom-developed power management board offers overcurrent protection, weighs less, and is more compact than off-the-shelf components. The gripper features two palms, both with two fingers inspired by the Fin Ray® effect [16]. The fingers leverage their compliance to bend around target objects upon contact, creating reliable grasps regardless of object geometry. The fingers are 3D-printed with 85A thermoplastic polyurethane to offer the necessary flexibility and strength for effective manipulation.
@@ 3.2 Middleware
Our software design follows the typical autonomy pipeline, including perception, planning, and control routines. The logic flow is illustrated in Figure 2, and the architecture of the software components is shown in Figure 3. For low-latency communication, we use ROS 2 [17] instead of ROS 1 [18], which is still the predominant middleware in most current works [3, 7, 13, 19, 20]. Unlike ROS 1, ROS 2 supports real-time communication, which is highly advantageous for applications requiring aggressive flight maneuvers [3, 4, 5].
@@ 3.3 Planning, Perception, Control
[…]
For planning, a mission control module manages the overall mission logic, observes the current state, and detects and handles exception states or failure modes. Working closely with this module, a trajectory planner computes the subsequent flight maneuvers based on the current mission state and specifies reference trajectories to the Pixhawk flight controller. The Pixhawk 6C runs PX4 Autopilot [21], an open-source flight control system that communicates over MAVLink [22]. To control the gripper in parallel, a separate ROS 2 interface receives commands over the network and forwards them to an Arduino Nano via UART serial communication.
[…]
@@ 3.4 Self-Localization
Achieving robust motion estimation through visual observations remains a challenging problem. While different open-source methods exist that perform well in synthetic benchmarks [24, 25, 26, 27, 28], we find that most of these published methods suffer from unsatisfactory robustness across different real-world scenarios. Multiple problems arise: high sensitivity to the camera and IMU calibration values makes it challenging to get repeatable localization results due to minor differences in calibration. Other parameters are often fine-tuned for specific synthetic benchmarks, necessitating a lengthy search to make pipelines work across different real-world environments.
The modular design of our pipeline allows for seamless integration of various SLAM algorithms. We conducted a comprehensive comparison of several open-source options, including SVO [28], Kimera [27], Rovio [24], and RTAB-Map [29]. Our evaluation revealed that Spectacular AI [10] consistently provided the most robust and precise state estimates. Notably, the modular nature of our system enables researchers to easily substitute any alternative SLAM pipelines as baselines when developing new state estimation algorithms. Further comparison details can be found in the Appendix (Section 6.3).
@@ 3.5 Target Localization
We propose a novel target localization pipeline, which can be used for zero-shot grasping of different objects. In contrast to existing pipelines for aerial manipulation [5, 12], this approach requires neither CAD models nor a semantic class label of the target objects, which enables strong generalization for zero-shot grasping.
The pipeline is visualized in Figure 4. The process is initialized with input from the operator, who selects a point of the target object in a livestream of the onboard camera. To produce a segmentation mask across subsequent image frames, we use SAM2 [30]. Given the segmentation mask and its corresponding depth image, we obtain a partial point cloud of the object.
We select the grasping points [31] using the following strategy. First, we estimate the centroid of the point cloud and determine its base. We then duplicate the point cloud and rotate it by 180 degrees (assuming symmetry of the object) to estimate its front and back. Using this assumed symmetry, we determine a new centroid for the object. Finally, we determine a set of candidate points for grasping. We do so by taking the intersection of the point cloud and a cutting plane that goes through the estimated centroid and is normal to the longest axis.
@@ 4.1 Flight Experiments
Figure 5: We evaluate the accuracy of the Spectacular AI SLAM system against motion capture ground truth during a sample grasping mission. All coordinates are expressed in NED (north-east-down) convention. The two trajectories have been aligned using Umeyama alignment with EVO [32]. We observe a peak RPE of 0.042 during a fast descent towards the object, whereas the mean RPE is 0.028.
@@ 4.2 Grasping Experiments
[…]
@@ 4.3 Results
In Table 1, we present the success rates for nine experiments obtained in the indoor setting using onboard perception only. Across a total of 144 grasp attempts, our system exhibits an average grasping success rate of 85%, which falls in line with previous works that rely on external perception systems [3, 4, 12]. In contrast to these works, we show zero-shot generalization to different objects by eliminating the need for explicit object priors in our grasp planning algorithm.
@@
[…]
@@ 4.4 Limitations
Our grasp planning algorithm operates on partial object point clouds and assumes approximate symmetry of target objects. This assumption limits the grasp planning capabilities of the platform, which could be addressed using point cloud reconstruction techniques [33] Furthermore, long-term video segmentation and tracking is not possible with SAM2, as the amount of memory and compute time increases with each frame. While sufficient for short grasping missions, using only short-term video segmentation (less than 400 frames) could be a limitation for other usages.

### v2

@@ Abstract
Figure 1: Our autonomous aerial platform demonstrates its aerial manipulation capabilities with onboard-only perception.
@@ 1 Introduction
Aerial manipulation, which combines the agility of aerial vehicles with the functional capabilities of manipulators, poses a significant challenge in robotics. Traditional mobile manipulation platforms, typically composed of a mobile base and a manipulator, are designed to interact with their environment and perform various tasks. Aerial manipulators, using multicopters as their mobile base, offer significant advantages over ground-based platforms such as legged robots [Bellicoso2019ALMAA] and wheeled robots [Klemm2019AscentoAT], particularly in scenarios requiring high traversal speeds or navigation through impassable terrain, such as during floods or in debris-strewn areas.
In recent years, aerial manipulation has seen significant advances using soft robotic manipulators [Fishman2021DynamicGW, Appius2022RAPTORRA, Ubellacker2023AggressiveAG, Cheung2023AMP]. The compliant properties of soft manipulators relax the otherwise high constraints on precisely positioning the manipulator when grasping objects. Furthermore, soft manipulators also dampen the impact on the dynamics of the flying platform, which eases the burden on flight controllers during each grasp.
However, a significant limitation of existing aerial manipulation platforms is their dependence on external perception systems, such as motion capture systems [Appius2022RAPTORRA, Ubellacker2023AggressiveAG, Xu2023AerialSM], fiducial markers [Fishman2021DynamicGW, Bianchi2023AND], and VIVE trackers [Lieret2020ALL], to perform self-localization and target object localization. These systems simplify the research problem by providing nearly perfect pose information, but also present two core issues.
@@
To overcome these limitations, we propose a new platform that relies exclusively on onboard perception sensors for aerial manipulation. For self-localization, we use an existing simultaneous localization and mapping (SLAM) system [spectacular2024], allowing the platform to estimate its state using visual observations alone. We implement a learning-based pipeline to localize target objects through RGB-D segmentation and point cloud analysis to determine target grasp locations.
@@ 2 Related Work
The integration of aerial systems with grasping mechanisms has gained considerable attention in recent years [Zhong2024PrototypeMA]. Research has demonstrated the potential of aerial grasping using soft grippers, which adapt to the shapes of objects due to their passive compliance [Fishman2021DynamicGW]. Further advances include rapid grasping platforms [Appius2022RAPTORRA] and efforts to achieve more autonomous grasping without relying on fiducial markers [Bauer2022AutonomousMR].
However, many existing systems still depend on motion capture systems to provide the precise location of the aerial platform or the object. For example, several works have focused on improving flying manipulators with stiffness control to improve object interaction dynamics [Sewalt2024EnhancingUC] and developing lightweight, low activation force grippers designed specifically for aerial grasping applications [Kremer2022ALU]. In addition, Wu et. al. [Wu2023RingRotorAN] has investigated optimizing quadcopter designs to enhance aerial grasping performance.
Despite these advancements, the reliance on motion capture systems for self-localization and object priors for target localization remains a common theme. More recent work has demonstrated successful aggressive aerial grasping using onboard perception by employing visual-inertial odometry for self-localization and CAD models for object recognition [Ubellacker2023AggressiveAG]. Our approach differs from this by leveraging a learning-based target localization system for zero-shot video segmentation integrated with a SLAM pipeline. This method eliminates the need for prior assumptions regarding object geometry or the environment, enhancing the flexibility and applicability of aerial grasping systems in unstructured settings.
@@ 3.1.1 Platform
The platform has an x-shaped configuration optimized for minimized and symmetric inertia in roll and pitch. Our multi-level design requires a vertical assembly: gripper components are located at the bottom, electronic circuit boards and flight controllers occupy the middle level, and the onboard computer and gripper controller are mounted above. This layout simplifies access to components that require frequent adjustments. Acrylic plates provide visual access while protecting the internal components. Electronic speed controllers (ESCs) are integrated within carbon profile arms, and frame connections are made of 3D-printed polylactic acid. The takeoff weight of the platform, including the battery, is $2479\text{\,}\mathrm{g}$, resulting in a payload capacity of approximately $500\text{\,}\mathrm{g}$.
A custom-developed power management board offers overcurrent protection, weighs less, and is more compact than off-the-shelf components. The gripper features two palms, both with two fingers inspired by the Fin Ray® effect [Crooks2016FinRE]. The fingers leverage their compliance to bend around target objects upon contact, creating reliable grasps regardless of object geometry. The fingers are 3D-printed with 85A thermoplastic polyurethane to offer the necessary flexibility and strength for effective manipulation.
@@ 3.2 Middleware
Our software design follows the typical autonomy pipeline, including perception, planning, and control routines. The logic flow is illustrated in Figure 2, and the architecture of the software components is shown in Figure 3. For low-latency communication, we use ROS 2 [Macenski2022RobotOS] instead of ROS 1 [Quigley2009ROSAO], which is still the predominant middleware in most current works [Fishman2021DynamicGW, Xu2023AerialSM, Sewalt2024EnhancingUC, RamnSoria2020GraspPA, Nedungadi2019DesignOA]. Unlike ROS 1, ROS 2 supports real-time communication, which is highly advantageous for applications requiring aggressive flight maneuvers [Fishman2021DynamicGW, Appius2022RAPTORRA, Ubellacker2023AggressiveAG].
@@ 3.3 Planning, Perception, Control
Our visual perception system uses an OAK-D S2 camera, which provides aligned grayscale and depth images at $30\text{\,}\mathrm{Hz}$ to both the SLAM pipeline and the target localization module. The SLAM pipeline also incorporates IMU data streamed at $200\text{\,}\mathrm{Hz}$. Using a single camera for both SLAM and downstream perception, we reduce overall weight and complexity compared to setups that use separate cameras for visual-inertial odometry (VIO) or SLAM and downstream perception [Ubellacker2023AggressiveAG].
For planning, a mission control module manages the overall mission logic, observes the current state, and detects and handles exception states or failure modes. Working closely with this module, a trajectory planner computes the subsequent flight maneuvers based on the current mission state and specifies reference trajectories to the Pixhawk flight controller. The Pixhawk 6C runs PX4 Autopilot [PX4], an open-source flight control system that communicates over MAVLink [MAVLink]. To control the gripper in parallel, a separate ROS 2 interface receives commands over the network and forwards them to an Arduino Nano via UART serial communication.
During testing, we frequently employ a VICON motion capture system to obtain ground-truth pose information on the quadcopter and the object. These ground-truth estimates are also used to test various modules in isolation. To access the VICON system measurements and publish them to the ROS 2 network, we employ an open-source repository from OPT4SMART [opt4smart2024]. The VICON system is configured to provide measurements at $200\text{\,}\mathrm{Hz}$.
@@ 3.4 Self-Localization
Achieving robust motion estimation through visual observations remains a challenging problem. While different open-source methods exist that perform well in synthetic benchmarks [Bloesch2017IteratedEK, MurArtal2015ORBSLAMAV, Qin2019AGO, Rosinol2019KimeraAO, Forster2017SVOSV], we find that most of these published methods suffer from unsatisfactory robustness across different real-world scenarios. Multiple problems arise: high sensitivity to the camera and IMU calibration values makes it challenging to get repeatable localization results due to minor differences in calibration. Other parameters are often fine-tuned for specific synthetic benchmarks, necessitating a lengthy search to make pipelines work across different real-world environments.
The modular design of our pipeline allows for seamless integration of various SLAM algorithms. We conducted a comprehensive comparison of several open-source options, including SVO [Forster2017SVOSV], Kimera [Rosinol2019KimeraAO], Rovio [Bloesch2017IteratedEK], and RTAB-Map [labbe2019rtab]. Our evaluation revealed that Spectacular AI [spectacular2024] consistently provided the most robust and precise state estimates. Notably, the modular nature of our system enables researchers to easily substitute any alternative SLAM pipelines as baselines when developing new state estimation algorithms. Further comparison details can be found in the Appendix (Section 6.3).
@@ 3.5 Target Localization
We propose a novel target localization pipeline, which can be used for zero-shot grasping of different objects. In contrast to existing pipelines for aerial manipulation [Ubellacker2023AggressiveAG, Bauer2022AutonomousMR], this approach requires neither CAD models nor a semantic class label of the target objects, which enables strong generalization for zero-shot grasping.
The pipeline is visualized in Figure 4. The process is initialized with input from the operator, who selects a point of the target object in a livestream of the onboard camera. To produce a segmentation mask across subsequent image frames, we use SAM2 [ravi2024sam2]. Given the segmentation mask and its corresponding depth image, we obtain a partial point cloud of the object.
We select the grasping points [ZapataImpata2019FastGC] using the following strategy. First, we estimate the centroid of the point cloud and determine its base. We then duplicate the point cloud and rotate it by 180 degrees (assuming symmetry of the object) to estimate its front and back. Using this assumed symmetry, we determine a new centroid for the object. Finally, we determine a set of candidate points for grasping. We do so by taking the intersection of the point cloud and a cutting plane that goes through the estimated centroid and is normal to the longest axis.
@@ 4.1 Flight Experiments
Figure 5: We evaluate the accuracy of the Spectacular AI SLAM system against motion capture ground truth during a sample grasping mission. All coordinates are expressed in NED (north-east-down) convention. The two trajectories have been aligned using Umeyama alignment with EVO [grupp2017evo]. We observe a peak RPE of 0.042 during a fast descent towards the object, whereas the mean RPE is 0.028.
@@ 4.2 Grasping Experiments
Figure 7: The eight objects we use for grasping experiments shown in Table 1. From left to right: plush toy, styrofoam object, cardboard box, ramen cup, bottle, kitchen roll, pouch, ball. The objects differ distinctly in size, weight, shape and surface friction. The bar denotes a length of $30\text{\,}\mathrm{cm}$.
@@ 4.3 Results
In Table 1, we present the success rates for nine experiments obtained in the indoor setting using onboard perception only. Across a total of 144 grasp attempts, our system exhibits an average grasping success rate of 85%, which falls in line with previous works that rely on external perception systems [Fishman2021DynamicGW, Appius2022RAPTORRA, Bauer2022AutonomousMR]. In contrast to these works, we show zero-shot generalization to different objects by eliminating the need for explicit object priors in our grasp planning algorithm.
@@
Object Success Rate Attempts Weight [ $\text{\,}\mathrm{g}$] Dimensions [ $\text{\,}\mathrm{cm}$]
@@ 4.4 Limitations
Our grasp planning algorithm operates on partial object point clouds and assumes approximate symmetry of target objects. This assumption limits the grasp planning capabilities of the platform, which could be addressed using point cloud reconstruction techniques [yuan2018pcn] Furthermore, long-term video segmentation and tracking is not possible with SAM2, as the amount of memory and compute time increases with each frame. While sufficient for short grasping missions, using only short-term video segmentation (less than 400 frames) could be a limitation for other usages.
@@ Acknowledgments
6 Appendix
6.1 Parts List
Additional resources for replicating our hardware platform are available in our public repository at https://github.com/srl-ethz/osprey. All files for sketches, 3D-printed linkages, laser-cut plates and an assembly guide can be found there. Given an RC Control unit, a workshop, and miscellaneous small parts such as screws, the total cost should be around \$ 1274. See Table 2 for the item list.
Count Name Price per Unit
1 Odroid H3+ \$ 165
1 NVME SSD \$ 50
2 16GB DDR4 RAM \$ 49
1 OAK-D \$ 250
1 Pixhawk 6C \$ 140
4 Motor 2216-920KV \$ 20
4 Propeller 1045 \$ 6
4 20A ESC \$ 14
1 LiPo 4300 mAh \$ 80
1 Arduino Nano \$ 20
1 Telemetry Transmitter \$ 60
1 Custom PCB \$ 40
1 3D printed parts \$ 30
4 Carbon Arms \$ 50
3 Acrylic Plates \$ 10
Total Price \$ 1274
Table 2: Total Price of the Platform
6.2 Power Management Board
(a) Top side of the custom PMB: a) on-off switches for power circuits b) DC/DC converters c) Arduino Nano d) power connector to onboard PC
(b) Bottom side of the custom PMB: a) servo connectors b) $5\text{\,}\mathrm{V}$ connectors c) lithium polymer battery input
Figure 8: 3D visualization of the custom power management board (PMB), designed to centralize and regulate power distribution for multiple voltage requirements across the platform.
Our custom PMB, shown in Figure 8, supplies power to the onboard computer ($19\text{\,}\mathrm{V}$ at up to $2\text{\,}\mathrm{A}$), the gripper’s servo motors ($6\text{\,}\mathrm{V}$ at up to $3\text{\,}\mathrm{A}$), and various onboard electronics ($5\text{\,}\mathrm{V}$ at up to $2\text{\,}\mathrm{A}$). The Arduino Nano is directly connected to the gripper’s servos to relay setpoints from the onboard computer.
6.3 Comparison of SLAM Algorithms
In the following, we go over different SLAM algorithms which we tested on our onboard computer. To support the use of ROS 1, we utilize the ros1_bridge package. We used each SLAM system with the provided default parameter set, adjusting it as needed to a reasonable degree. To this end, this overview is not exhaustive and further performance gains could likely be obtained by spending more time tuning the different algorithms. An in-depth quantitative overview and comparison of different SLAM algorithms can be explored in future work with a stronger focus on self-localization.
  • SVO Pro [Forster2017SVOSV]: very sensitive to camera calibration. Very little documentation. Compatible with ROS 1.
  • Kimera [Rosinol2019KimeraAO]: precise for slow maneuvers, but does not run fast enough for faster maneuvers without sacrificing accuracy. Compatible with ROS 1.
  • Rovio [Bloesch2017IteratedEK]: similar to Kimera, does not run fast enough to fly agile maneuvers without sacrificing accuracy. Compatible with ROS 1.
  • RTAB-MAP [labbe2019rtab]: similar to Kimera, does not run fast, relocalization capabilities are not robust enough for agile drone flight. Compatible with ROS 1.
  • Spectacular AI [spectacular2024]: offers out-of-the-box SLAM, which runs fast and offers good precision. Template wrappers are provided for ROS 1 and ROS 2. Closed-source core algorithm.
Other common tracking solutions include the Intel RealSense T265 camera, which offers out-of-the-box tracking, but is no longer produced and suffers from poor availability. Another off-the-shelf alternative is the VIO pipeline provided by ZED. However, the ZED SDK requires CUDA on the host computer, which greatly constrains the choice of onboard computers because it requires a CUDA-compatible graphical processing unit.